OpenStudy (anonymous):

find domain of f(x)=cos^-1(2-x/2)

OpenStudy (gorv):

u know the domain of cos inverse??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

domain is from -1 to 1

OpenStudy (gorv):

so what is inside should be inside this

OpenStudy (gorv):

\[-1 \le 2-\frac{ x }{ 2 } \le 1\]

OpenStudy (gorv):

subtract 2 from each now

OpenStudy (gorv):

n change the sign of inequality

OpenStudy (akonkel):

@danish071996